OpenStudy (anonymous):

In simple harmonic motion what is the logic behind omega squared equaling k/m?

OpenStudy (vincent-lyon.fr):

\(\omega_o\) is just an easy way to avoid writing \(\sqrt{\Large\frac km}\) in the solution \(x(t)=A \cos(\sqrt{\Large\frac km}t+\phi)\) which would be a bit tedious in the long term.